OpenStudy (anonymous):

Plz help! Ill give medal and fan! but i really need help... Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

A. (3, 1) B. (-1, 3) C. (-1, -3) D. (1, 3)

OpenStudy (anonymous):

@mathkid911 Plz help

OpenStudy (anonymous):

@Bee_see can u help?

OpenStudy (anonymous):

plug in the coordinates in each equation and see if they are equal

OpenStudy (anonymous):

so for example if (3,1) was the right answer you would plug in 3 for x and 1 for y and calculate if the two equations would be the same

OpenStudy (anonymous):

ohh ok let me check :)

OpenStudy (bee_see):

Ok...You want to get rid of a variable. You can get rid of y by subtracting y-y, right? 1y-1y=0.

OpenStudy (bee_see):

-2y-(-4x)=2x... 1-(-1)=2

OpenStudy (anonymous):

yes

OpenStudy (bee_see):

we are left with 2x=2....divide by 2. x=1

OpenStudy (bee_see):

put any of the two equations. insert x.I chose the first one. so...

OpenStudy (bee_see):

-2(1)+y=1

OpenStudy (bee_see):

-2+y=1

OpenStudy (bee_see):

add by 2...we get y=3.

OpenStudy (anonymous):

ok so B and C are out right?

OpenStudy (bee_see):

Check the variables! -2(1)+(3)=1 and -4(1)+3=-1

OpenStudy (anonymous):

So is it D?

OpenStudy (anonymous):

yes i got that part

OpenStudy (bee_see):

D

OpenStudy (anonymous):

@Bee_see is it D?

OpenStudy (bee_see):

Yes. Do you understand how I got the answer?

OpenStudy (anonymous):

yes thx for your help!

OpenStudy (anonymous):

can u help me with one more?

OpenStudy (bee_see):

Ok. Lets see if I can help.

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1

OpenStudy (anonymous):

x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (bee_see):

Ok.Which variable do you want to eliminate?

OpenStudy (anonymous):

idk

OpenStudy (bee_see):

ok. Let me work on it then.

OpenStudy (anonymous):

ok thx

OpenStudy (bee_see):

infinite solutions

OpenStudy (anonymous):

D? If so can u explain how u got it plz

OpenStudy (anonymous):

@Bee_see r u there?

OpenStudy (bee_see):

ah. sorry. no solution.

OpenStudy (anonymous):

thx

OpenStudy (bee_see):